# Grassmann–Plücker functions for orthogonal matroids

### v1

@@ 1 Introduction
The equivalence between the first and third was proven in [14], and the second is the new cryptomorphism. Our result is even new in the realizable case. Indeed, when $F$ is the complex field, we have the following result. For the formal statement, see §3.2.
@@
(Cor. 3.7) Let $p^{\prime}:\mathrm{OG}(n,2n)\to\mathbb{P}^{2^{n}+n(n-1)2^{n-2}-1}(\mathbb{C})$ be the restricted Plücker embedding; that is, the image of $V$ is the restriction of the Plücker coordinates $p(V)$ to the coordinates $I\in\binom{[n]\cup[n]^{*}}{n}$ such that $I$ contains at most one $\{i,i^{*}\}$ with $1\leq i\leq n$. Then the image of each connected component of $\mathrm{OG}(n,2n)$ under $p^{\prime}$ is cut out by the (restricted) Plücker relations along with certain linear relations.
We also present analogous equivalences for weak orthogonal $F$-matroids. As noted in [1], there are (at least) two natural notions of matroids over tracts, called weak and strong $F$-matroids, and similar situations occur in the study of orthogonal matroids [14]. The importance of weak $F$-matroids dates back to Tutte’s excluded-minor characterization of binary and regular matroids [22], [^1] Via the Tutte coordinatization theorem, one can effectively construct a representation of a matroid; see [18, §2.3]. Roughly speaking, it asserts that representability can be detected by only considering the modular pairs of circuits (or, dually, of hyperplanes). In the terminology of [1], Tutte’s theorem can be stated as follows: if $R$ is an integral domain, every weak $R$-matroid is strong. and is connected to the recent study of foundations of matroids [5, 4] and orthogonal matroids [15].
@@
Besides Cayley’s identity, our work is also inspired by the result of the second author [16] that develops the Baker–Bowler theory for type C which generalizes the embedding of the Lagrangian Grassmannian via principal and almost-principal minors of symmetric matrices described in [6].
@@ 2.2 Orthogonal matroids over tracts
Let $A$ be an $n$-by-$n$ skew-symmetric matrix over a field $K$. Then $\mathcal{B}=\{([n]\setminus X)\cup X^{*}\mid A_{X}\text{ is nonsingular}\}$ is the basis set of an orthogonal matroid, denoted $M(A)$, on $n]\cup[n]^{*}$, where $A_{X}$ is the principal submatrix of $A$ with rows and columns indexed by $X\subseteq[n]$. Given an orthogonal matroid $M=([n]\cup[n]^{*},\mathcal{B})$ and a set $R\subseteq[n]$, a pair $M\triangle R:=([n]\cup[n]^{*},\mathcal{B}^{\prime})$ is an orthogonal matroid, where $\mathcal{B}^{\prime}=\{B\triangle(R\cup R^{*})\mid B\in\mathcal{B}\}$. An orthogonal matroid $M$ is representable over $K$ if it is equal to $M(A)\triangle R$ for some matrix $A$ over $K$ and some set $R$. Bouchet [11] showed that a matroid is representable over $K$ if and only if its lift is representable over $K$.
@@
The space $\mathrm{OG}(n,2n)$ is embedded into $\mathbb{P}^{2^{n}-1}(\mathbb{C})$ via the Wick embedding $w$. For simplicity, we assume that $V$ is the row space of an $n$-by-$2n$ matrix $(I_{n}\mid A)$ with columns indexed by $1,2,\dots,n,1^{*},2^{*},\dots,n^{*}$ in this order. Then $A$ is skew-symmetric. The Wick coordinates $w(V)\in\mathbb{P}^{2^{n}-1}(\mathbb{C})$ are defined as
$w(V)(T)=\mathrm{pf}(A_{T\setminus[n]})\text{ for }T\in\mathcal{T}_{n},$
where $A_{T\setminus[n]}$ is the principal submatrix of $A$ with columns indexed by $T\setminus[n]$.
@@ 3.1 The new cryptomorphism
    There is $\sigma\in\{0,1\}$ such that $\varphi(B)=0$ for any $B\in\mathcal{T}_{n}^{1-\sigma}$.
@@ 3.2 Restricted Grassmann–Plücker functions over the complex field
$p(V)(I):=\det(A_{I})\text{ for $n$-subsets }I\text{ of }[n]\cup[n]^{*},$
where $A$ is an $n$-by-$2n$ matrix whose row space is $V$ and whose columns are indexed by $1<1^{*}<2<2^{*}<\dots<n<n^{*}$ in order, and $A_{I}$ is the submatrix of $A$ with columns indexed by $I$. In the case that $A_{[n]}$ is the identity matrix (so, $V\in OG^{+}(n,2n)$), the submatrix $A_{[n]^{*}}$ is skew-symmetric and thus for all $B\in\mathcal{T}_{n}^{0}$ and distinct $i,j\in[n]$,
@@
Let $A$ and $A^{\prime}$ be the following 22-by-44 matrices:
@@
where $a\in\mathbb{C}$ and the columns are indexed by $1<1^{*}<2<2^{*}$ in order. We denote by $V$ and $V^{\prime}$ the row spaces of $A$ and $A^{\prime}$, respectively. Then $V\in\mathrm{OG}^{+}(2,4)$ and $V^{\prime}\in\mathrm{OG}^{-}(2,4)$. If we take $B=\{1,2\}\in\mathcal{T}_{n}^{0}$, $i=1$, and $j=2$, then we have:
$p(V)(B\setminus\{1,1^{*}\}\cup\{2,2^{*}\})$ $=\det(A_{\{2,2^{*}\}})=-a,$
$p(V)(B\cup\{1,1^{*}\}\setminus\{2,2^{*}\})$ $=\det(A_{\{1,1^{*}\}})=-a.$
@@
$p(V^{\prime})(B^{\prime}\setminus\{1,1^{*}\}\cup\{2,2^{*}\})=\det(A^{\prime}_{\{2,2^{*}\}})$ $=-a,$
$p(V^{\prime})(B^{\prime}\cup\{1,1^{*}\}\setminus\{2,2^{*}\})=\det(A^{\prime}_{\{1,1^{*}\}})$ $=a.$
@@
(Cor. 1.2) Let $p^{\prime}:\mathrm{OG}(n,2n)\to\mathbb{P}^{2^{n}+n(n-1)2^{n-2}-1}(\mathbb{C})$ be the restricted Plücker embedding; that is, the image of $V$ is the restriction of the Plücker coordinates $p(V)$ to coordinates $I$ in $\mathcal{T}_{n}\cup\mathcal{A}_{n}$. Then the image of $\mathrm{OG}^{+}(n,2n)$ under $p^{\prime}$ is cut out by the restricted Plücker relations along with the linear relations: for each $B\in\mathcal{T}_{n}^{0}$ and distinct $i,j\in[n]$,
@@ 3.4.1 Wick functions to restricted GP functions
When $F$ is a field, the map $\Phi_{\mathrm{W}\to\mathrm{GP}}$ is derived from Cayley’s identities [17, Equation (6.3)]: for an $n\times n$ skew-symmetric matrix $A$,
$\det(A)=\mathrm{pf}(A)^{2}\ \text{ and }\ \det(A_{j}^{i})=\begin{cases}(-1)^{\mathbf{1}(i>j)}\mathrm{pf}(A_{ij}^{ij})\mathrm{pf}(A)&\text{$n$ is even},\\ \mathrm{pf}(A_{i}^{i})\mathrm{pf}(A_{j}^{j})&\text{$n$ is odd},\end{cases}$
where $A_{j}^{i}$ is the submatrix obtained by removing the $i$-th column and $j$-th row. See Appendix A for more discussions.
@@ 5.2 Almost-principal minors and Pfaffian positivity
For a square matrix $A$, an almost-principal submatrix $A[I\cup\{i\},I\cup\{j\}]$ is called top-right if $i<j$.
@@
Let $A$ be an $n$-by-$n$ skew-symmetric matrix with real entries. Then the following are equivalent:
@@
Suppose $A=(a_{ij})$ is a skew-symmetric matrix whose odd-sized top-right almost-principal minors are nonnegative. We show that $\mathrm{pf}(A_{I})\geq 0$ for all even-sized subsets $I\subseteq[n]$. We proceed by induction on |I|/2{|I|}/{2}. The base case $I=\emptyset$ is trivial, so we assume that $|I|>0$. We may assume $\mathrm{pf}(A_{I})\neq 0$. Denote $I=\{i_{1}<i_{2}<\dots<i_{2k}\}$. Because of the recursive definition of Pfaffians, we have
$\mathrm{pf}(A_{I})=\sum_{j=2}^{2k}(-1)^{j}a_{i_{1}i_{j}}\mathrm{pf}(A_{I\setminus\{i_{1},i_{j}\}}).$
Then $\mathrm{pf}(A\setminus\{i_{1},i_{j}\})$ is nonzero for some $j$, which is indeed positive by induction hypothesis. By Cayley’s Pfaffian identity, we conclude that
$\mathrm{pf}(A_{I})=\frac{\det(A[I\setminus\{i_{j}\},I\setminus\{i_{1}\}])}{\mathrm{pf}(A_{I\setminus\{i_{1},i_{j}\}})}>0.\qed$
@@ Appendix A $\Phi_{\mathrm{W}\to\mathrm{GP}}$ in the realizable case
We give an intuition for the map $\Phi_{\mathrm{W}\to\mathrm{GP}}$, defined in §3.4.1, by looking at the realizable case.
Let $A$ be an $n$-by-$n$ skew-symmetric matrix over a field $K$. Then, the corresponding Wick $K$-function $\psi:\mathcal{T}_{n}\to K$ is defined by
$\psi(([n]\setminus I)\cup I^{*}):=\mathrm{pf}(A[I,I])$
for $I\subseteq[n]$, where $A[R,C]$ denotes the submatrix with rows indexed by $R$ and columns indexed by $C$.
Let $A^{\prime}$ be the $n$-by-$2n$ matrix with columns indexed by $1,1^{*},2,2^{*},\dots,n,n^{*}$ in order such that the square submatrix with columns $1,\dots,n$ is the identity matrix and the square submatrix with columns $1^{*},\dots,n^{*}$ is $A$. The corresponding restricted GP $K$-function $\varphi:\mathcal{T}_{n}\cup\mathcal{A}_{n}\to K$ is defined by
$\varphi(B):=\det(A^{\prime}[B])$
for $B\in\mathcal{T}_{n}\cup\mathcal{A}_{n}$, where $A^{\prime}[B]$ denotes the square submatrix with columns indexed by $B$. Note that $\sigma=0$ in this case, due to $\varphi([n])=1\neq 0$.
@@
$\varphi(([n]\setminus I)\cup I^{*})=\det(A^{\prime}[([n]\setminus I)\cup I^{*}])$ $=\det(A[I,I])$
$=\mathrm{pf}(A[I,I])^{2}=\psi(([n]\setminus I)\cup I^{*})^{2}.$
@@
The almost-transversal case (1) is more involved with respect to the sign, which eventually comes from the second part of Cayley’s Pfaffian identity:
$\det(A[Ii,Ij])=\begin{cases}(-1)^{\mathbf{1}(i>j)}\mathrm{pf}(A[I,I])\mathrm{pf}(A[Iij,Iij])&\text{$|I|$ is even},\\ \mathrm{pf}(A[Ii,Ii])\mathrm{pf}(A[Ij,Ij])&\text{$|I|$ is odd},\end{cases}$
@@
Here, $|X>i|$ denotes the number of elements $x\in X$ with $x>i$. The number $m$ comes from the cofactor expansions along the columns indexed by $n]\setminus Ii$. Then, one can derive (1) by computing the signs as follows.
@@
$\det(A[Ii,Ij])=(-1)^{\mathbf{1}(i>j)}\mathrm{pf}(A[I,I])\mathrm{pf}(A[Iij,Iij]).$
@@
$\varphi(([n]\setminus Ii)\cup(Ij)^{*})$ $=(-1)^{m}\det(A[Ii,Ij])$
$=(-1)^{m+\mathbf{1}(i>j)}\mathrm{pf}(A[I,I])\mathrm{pf}(A[Iij,Iij])$
@@
$\det(A[Ii,Ij])=\mathrm{pf}(A[Ii,Ii])\mathrm{pf}(A[Ij,Ij]),$
@@
$\varphi(([n]\setminus Ii)\cup(Ij)^{*})$ $=(-1)^{m}\det(A[Ii,Ij])$
$=(-1)^{m}\mathrm{pf}(A[Ii,Ii])\mathrm{pf}(A[Ij,Ij])$

### v2

@@ 1 Introduction
1.1 Main results
Our first main result is as follows.
@@ 1.1 Main results
The equivalence between the first and third was proven in [14], and the second is the new cryptomorphism. Our result is even new in the realizable case. Indeed, when $F$ is the complex field, we have the following result. For the formal statement, see §3.2. For an $n$-element subset $I$ of $n]\cup[n]^{*}:=\{1,2,\dots,n,1^{*},2^{*},\dots,n^{*}\}$, we call $I$ a transversal if it contains exactly one of $i$ and $i^{*}$ for each $1\leq i\leq n$, and call $I$ an almost-transversal if it contains exactly one pair of the form $\{i,i^{*}\}$. Also, recall that the orthogonal Grassmannian $\mathrm{OG}(n,2n)$ is the set of all $n$-dimensional isotropic subspaces of $\mathbb{C}^{[n]\cup[n]^{*}}$ with respect to a symmetric non-degenerate bilinear form.
@@
(Cor. 3.7) Let $p^{\prime}:\mathrm{OG}(n,2n)\to\mathbb{P}^{2^{n}+n(n-1)2^{n-2}-1}(\mathbb{C})$ be the restricted Plücker embedding; that is, the image of $V\in\mathrm{OG}(n,2n)$ is the Plücker coordinates indexed by transversals and almost-transversals. Then the image of each connected component of $\mathrm{OG}(n,2n)$ under $p^{\prime}$ is cut out by the restricted Plücker relations along with certain linear relations.
We also present analogous equivalences for weak orthogonal $F$-matroids. As noted in [1], there are (at least) two natural notions of matroids over tracts, called weak and strong $F$-matroids, and similar situations occur in the study of orthogonal matroids [14]. The study of weak $F$-matroids has a long history, predating the modern terminology. When $K$ is a field, a weak $K$-matroid is exactly a vector that satisfies all the $3$-term Plücker relations and whose support forms a matroid. It is a folklore fact that such a vector satisfies all the Plücker relations; see [3] for a proof. In his paper [21], Tutte studied chain-group representations of matroids and used this notion to find excluded-minors of binary and regular matroids. For any partial field $F$, the chain-group representations are (strong) $F$-circuit sets of matroids in the sense of [1], which are cryptomorphic to (strong) $F$-matroids. A key step in Tutte’s paper is to prove that weak $F$-circuit sets of matroids are in fact strong for any partial field $F$. More recently, weak $F$-matroids have appeared in the study of foundations of matroids [5, 4] and orthogonal matroids [15].
@@
1.2 Connections with other works.
An analogous result of Corollary 1.2 for the Lagrangian Grassmannian, which is of type C, was proved by Boege et al. [6]. They embedded the Lagrangian Grassmannian into the projective space through the restricted Plücker embedding described in Corollary 1.2, and showed that the image of the embedding is cut out by certain quadratic equations (along with certain linear relations). Based on their work, the second author [16] simplified their list of quadratic relations to the restricted Plücker relations. Our results are inspired by [6] and [16], but for the orthogonal Grassmannian which is of type D.
The readers who are familiar with [16] might notice that orthogonal matroids are a subclass of the antisymmetric matroids defined in [16], and the cryptomorphisms for antisymmetric $F$-matroid for a tract $F$ have been studied in the same paper. However, orthogonal $F$-matroids are not a subclass of antisymmetric $F$-matroids because one mimics the orthogonal Grassmannian and the other mimics Lagrangian Grassmannian. For a precise discussion, see Remark 3.14.
Our main contribution is to use the idea of restricted Plücker relations to study orthogonal $F$-matroids, which have previously been studied using Wick relations. We finally remark that it is difficult to find a type C analogue of Wick relations, because there is no known notion of Pfaffians for symmetric matrices.
1.3 Organization of the paper
@@ 2.2 Orthogonal matroids over tracts
Let $\mathbf{A}$ be an $n$-by-$n$ skew-symmetric matrix over a field $K$. Then $\mathcal{B}=\{([n]\setminus X)\cup X^{*}\mid\mathbf{A}_{X}\text{ is nonsingular}\}$ is the basis set of an orthogonal matroid, denoted $M(\mathbf{A})$, on $n]\cup[n]^{*}$, where $\mathbf{A}_{X}$ is the principal submatrix of $\mathbf{A}$ with rows and columns indexed by $X\subseteq[n]$. Given an orthogonal matroid $M=([n]\cup[n]^{*},\mathcal{B})$ and a set $R\subseteq[n]$, a pair $M\triangle R:=([n]\cup[n]^{*},\mathcal{B}^{\prime})$ is an orthogonal matroid, where $\mathcal{B}^{\prime}=\{B\triangle(R\cup R^{*})\mid B\in\mathcal{B}\}$. An orthogonal matroid $M$ is representable over $K$ if it is equal to $M(\mathbf{A})\triangle R$ for some matrix $\mathbf{A}$ over $K$ and some set $R$. Bouchet [11] showed that a matroid is representable over $K$ if and only if its lift is representable over $K$.
@@
The space $\mathrm{OG}(n,2n)$ is embedded into $\mathbb{P}^{2^{n}-1}(\mathbb{C})$ via the Wick embedding $w$. For simplicity, we assume that $V$ is the row space of an $n$-by-$2n$ matrix $(\mathbf{I}_{n}\mid\mathbf{A})$ with columns indexed by $1,2,\dots,n,1^{*},2^{*},\dots,n^{*}$ in this order. Then $\mathbf{A}$ is skew-symmetric. The Wick coordinates $w(V)\in\mathbb{P}^{2^{n}-1}(\mathbb{C})$ are defined as
$w(V)(T)=\mathrm{pf}(\mathbf{A}_{T\setminus[n]})\text{ for }T\in\mathcal{T}_{n},$
where $\mathbf{A}_{T\setminus[n]}$ is the principal submatrix of $\mathbf{A}$ with columns indexed by $T\setminus[n]$.
@@ 3.1 The new cryptomorphism
    There is (a unique) $\sigma\in\{0,1\}$ such that $\varphi(B)=0$ for any $B\in\mathcal{T}_{n}^{1-\sigma}$.
@@ 3.2 Restricted Grassmann–Plücker functions over the complex field
$p(V)(I):=\det(\mathbf{A}_{I})\text{ for $n$-subsets }I\text{ of }[n]\cup[n]^{*},$
where $\mathbf{A}$ is an $n$-by-$2n$ matrix whose row space is $V$ and whose columns are indexed by $1<1^{*}<2<2^{*}<\dots<n<n^{*}$ in order, and $\mathbf{A}_{I}$ is the submatrix of $\mathbf{A}$ with columns indexed by $I$. In the case that $\mathbf{A}_{[n]}$ is the identity matrix (so, $V\in OG^{+}(n,2n)$), the submatrix $\mathbf{A}_{[n]^{*}}$ is skew-symmetric and thus for all $B\in\mathcal{T}_{n}^{0}$ and distinct $i,j\in[n]$,
@@
Let $\mathbf{A}$ and $\mathbf{A}^{\prime}$ be the following 22-by-44 matrices:
@@
where $a\in\mathbb{C}$ and the columns are indexed by $1<1^{*}<2<2^{*}$ in order. We denote by $V$ and $V^{\prime}$ the row spaces of $\mathbf{A}$ and $\mathbf{A}^{\prime}$, respectively. Then $V\in\mathrm{OG}^{+}(2,4)$ and $V^{\prime}\in\mathrm{OG}^{-}(2,4)$. If we take $B=\{1,2\}\in\mathcal{T}_{n}^{0}$, $i=1$, and $j=2$, then we have:
$p(V)(B\setminus\{1,1^{*}\}\cup\{2,2^{*}\})$ $=\det(\mathbf{A}_{\{2,2^{*}\}})=-a,$
$p(V)(B\cup\{1,1^{*}\}\setminus\{2,2^{*}\})$ $=\det(\mathbf{A}_{\{1,1^{*}\}})=-a.$
@@
$p(V^{\prime})(B^{\prime}\setminus\{1,1^{*}\}\cup\{2,2^{*}\})=\det(\mathbf{A}^{\prime}_{\{2,2^{*}\}})$ $=-a,$
$p(V^{\prime})(B^{\prime}\cup\{1,1^{*}\}\setminus\{2,2^{*}\})=\det(\mathbf{A}^{\prime}_{\{1,1^{*}\}})$ $=a.$
@@
(Cor. 1.2) Let $p^{\prime}:\mathrm{OG}(n,2n)\to\mathbb{P}^{2^{n}+n(n-1)2^{n-2}-1}(\mathbb{C})$ be the restricted Plücker embedding defined by
$p^{\prime}(V)=(p(V)(I)\colon I\in\mathcal{T}_{n}\cup\mathcal{A}_{n}).$
Then the image of $\mathrm{OG}^{+}(n,2n)$ under $p^{\prime}$ is cut out by the restricted Plücker relations along with the linear relations: for each $B\in\mathcal{T}_{n}^{0}$ and distinct $i,j\in[n]$,
@@ 3.4.1 Wick functions to restricted GP functions
When $F$ is a field, the map $\Phi_{\mathrm{W}\to\mathrm{GP}}$ is derived from Cayley’s identities [17, Equation (6.3)]: for an $n\times n$ skew-symmetric matrix $\mathbf{A}$,
$\det(\mathbf{A})=\mathrm{pf}(\mathbf{A})^{2}\ \text{ and }\ \det(\mathbf{A}_{j}^{i})=\begin{cases}(-1)^{\mathbf{1}(i>j)}\mathrm{pf}(\mathbf{A}_{ij}^{ij})\mathrm{pf}(\mathbf{A})&\text{$n$ is even},\\ \mathrm{pf}(\mathbf{A}_{i}^{i})\mathrm{pf}(\mathbf{A}_{j}^{j})&\text{$n$ is odd},\end{cases}$
where $\mathbf{A}_{j}^{i}$ is the submatrix obtained by removing the $i$-th column and $j$-th row. See Appendix A for more discussions.
@@ 5.1 Connection to enveloping matroids
Note that when $F$ is the Krasner hyperfield, the question is exactly asking whether any orthogonal matroid admits an enveloping matroid. When $F$ is a field, from the Wick function $\psi$, we may construct a skew-symmetric matrix $\mathbf{A}$ whose Pfaffians are the values of $\psi$, and the $n$-by-$2n$ matrix $(\mathbf{I}_{n}\mid\mathbf{A})$ represents an $F$-matroid having the desired property.
@@ 5.2 Almost-principal minors and Pfaffian positivity
For a square matrix $\mathbf{A}$, an almost-principal submatrix $\mathbf{A}[I\cup\{i\},I\cup\{j\}]$ is called top-right if $i<j$.
@@
Let $\mathbf{A}$ be an $n$-by-$n$ skew-symmetric matrix with real entries. Then the following are equivalent:
@@
Suppose $\mathbf{A}=(a_{ij})$ is a skew-symmetric matrix whose odd-sized top-right almost-principal minors are nonnegative. We show that $\mathrm{pf}(\mathbf{A}_{I})\geq 0$ for all even-sized subsets $I\subseteq[n]$. We proceed by induction on |I|/2{|I|}/{2}. The base case $I=\emptyset$ is trivial, so we assume that $|I|>0$. We may assume $\mathrm{pf}(\mathbf{A}_{I})\neq 0$. Denote $I=\{i_{1}<i_{2}<\dots<i_{2k}\}$. Because of the recursive definition of Pfaffians, we have
$\mathrm{pf}(\mathbf{A}_{I})=\sum_{j=2}^{2k}(-1)^{j}a_{i_{1}i_{j}}\mathrm{pf}(\mathbf{A}_{I\setminus\{i_{1},i_{j}\}}).$
Then $\mathrm{pf}(\mathbf{A}\setminus\{i_{1},i_{j}\})$ is nonzero for some $j$, which is indeed positive by induction hypothesis. By Cayley’s Pfaffian identity, we conclude that
$\mathrm{pf}(\mathbf{A}_{I})=\frac{\det(\mathbf{A}[I\setminus\{i_{j}\},I\setminus\{i_{1}\}])}{\mathrm{pf}(\mathbf{A}_{I\setminus\{i_{1},i_{j}\}})}>0.\qed$
@@ Appendix A $\Phi_{\mathrm{W}\to\mathrm{GP}}$ in the realizable case
In §3.4.1, we construct the GP function $\varphi$ from a Wick function $\psi$, and define $\Phi_{\mathrm{W}\to\mathrm{GP}}([\psi]):=[\varphi]$. The formulas involved in the construction are complicated, especially the sign in (1). In this appendix, we justify our construction by showing that it is correct in the realizable case.
Let $\mathbf{A}$ be an $n$-by-$n$ skew-symmetric matrix over a field $K$. Then, the corresponding Wick $K$-function $\psi:\mathcal{T}_{n}\to K$ is defined by
$\psi(([n]\setminus I)\cup I^{*}):=\mathrm{pf}(\mathbf{A}[I,I])$
for $I\subseteq[n]$, where $\mathbf{A}[R,C]$ denotes the submatrix with rows indexed by $R$ and columns indexed by $C$.
Let $\mathbf{A}^{\prime}$ be the $n$-by-$2n$ matrix with columns indexed by $1,1^{*},2,2^{*},\dots,n,n^{*}$ in order such that the square submatrix with columns $1,\dots,n$ is the identity matrix and the square submatrix with columns $1^{*},\dots,n^{*}$ is $\mathbf{A}$. We consider the restricted GP $K$-function $\varphi:\mathcal{T}_{n}\cup\mathcal{A}_{n}\to K$ defined by
$\varphi(B):=\det(\mathbf{A}^{\prime}[B])$
for $B\in\mathcal{T}_{n}\cup\mathcal{A}_{n}$, where $\mathbf{A}^{\prime}[B]$ denotes the square submatrix with columns indexed by $B$. Note that $\sigma=0$ in this case, due to $\varphi([n])=1\neq 0$.
Our target is to show that this $\varphi$ is exactly the one we defined in §3.4.1.
@@
$\varphi(([n]\setminus I)\cup I^{*})=\det(\mathbf{A}^{\prime}[([n]\setminus I)\cup I^{*}])$ $=\det(\mathbf{A}[I,I])$
$=\mathrm{pf}(\mathbf{A}[I,I])^{2}=\psi(([n]\setminus I)\cup I^{*})^{2}.$
@@
It remains to show the almost-transversal case (1). We will use the second part of Cayley’s Pfaffian identity:
$\det(\mathbf{A}[Ii,Ij])=\begin{cases}(-1)^{\mathbf{1}(i>j)}\mathrm{pf}(\mathbf{A}[I,I])\mathrm{pf}(\mathbf{A}[Iij,Iij])&\text{$|I|$ is even},\\ \mathrm{pf}(\mathbf{A}[Ii,Ii])\mathrm{pf}(\mathbf{A}[Ij,Ij])&\text{$|I|$ is odd},\end{cases}$
@@
Here, $|X>i|$ denotes the number of elements $x\in X$ with $x>i$. The number $m$ comes from the cofactor expansion along the columns indexed by $n]\setminus Ii$. Then, one can derive (1) by computing the signs as follows.
@@
$\det(\mathbf{A}[Ii,Ij])=(-1)^{\mathbf{1}(i>j)}\mathrm{pf}(\mathbf{A}[I,I])\mathrm{pf}(\mathbf{A}[Iij,Iij]).$
@@
$\varphi(([n]\setminus Ii)\cup(Ij)^{*})$ $=(-1)^{m}\det(\mathbf{A}[Ii,Ij])$
$=(-1)^{m+\mathbf{1}(i>j)}\mathrm{pf}(\mathbf{A}[I,I])\mathrm{pf}(\mathbf{A}[Iij,Iij])$
@@
$\det(\mathbf{A}[Ii,Ij])=\mathrm{pf}(\mathbf{A}[Ii,Ii])\mathrm{pf}(\mathbf{A}[Ij,Ij]),$
@@
$\varphi(([n]\setminus Ii)\cup(Ij)^{*})$ $=(-1)^{m}\det(\mathbf{A}[Ii,Ij])$
$=(-1)^{m}\mathrm{pf}(\mathbf{A}[Ii,Ii])\mathrm{pf}(\mathbf{A}[Ij,Ij])$